Fermi National Laboratory
 

Operation Summary for October 6, 2010 - October 8, 2010

 

Wednesday Oct. 6
The day shift began with Operations monitoring the Tevatron (TeV) store 8145 (79.50E30) and the Recycler stash (178.6E10), with theAntiproton Source (Pbar) stacking (20.6E10), with the Linac, Booster, and theMain Injector(MI) operating normally, with NuMI off for target work, with MiniBooNE taking beam, and with theCollider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8145 at 9:57 AM due to a vacuum leak on the MI-52 Lambertson. Operations put the TeV into HEP shot setup. Operations halted the Booster beam at 10:40 AM for an access to repair a RF station (BRF18). Operations established store 8146 at 11:06 AM with an average initial luminosity (AIL) of 228.05E30. Mechanical technicians accessed MI to investigate vacuum problems at MI-52. The techs discovered a leaky ion pump high-voltage feed-through and began repairs.

During the evening shift, the Mechanical techs completed their MI vacuum repairs at 7:10 PM. Operations established beam to Pbar and MiniBooNE at 7:45 PM. Operations reported downtime due to the following: LRF3, a bad MI60S CIA crate card, and a Booster East Anode Power Supply (EAPS) trip.

Thursday Oct. 7
The midnight shift began with Operations monitoring store 8146 (42.35E30) and the stash (124.4E10), with Pbar stacking (13.5E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8146 at 7 AM. Operators prepared Enclosure B and the D-Zero collision hall for access.

During the day shift, D-Zero personnel came out of access at 9:30 AM. Operations checked the TeV orbits and put the TeV into shot setup at 11:45 AM. The power amplifier (PA) for TRF1 faulted, prior to loading antiprotons, and needed to be replaced; RF personnel replaced the PA. Operations put the TeV back into shot setup at 2:40 PM and established store 8151 at 3:45 PM with an AIL of 318E30. The MiniBooNE Horn tripped off on a ground fault indication; MiniBooNE personnel investigated and increased the trip limit in lieu of a lengthy repair.

During the evening shift, subtle Preaccelerator high-voltage amplitude variations created problems for the low-energy Linac RF systems that inhibited beam intermittently. Linac personnel investigated and recalibrated the feedback loops for the various RF stations. Linac beam quality was unstable for less than half an hour.

Friday Oct. 8
The midnight shift began with Operations monitoring store 8151 (103.82E30) and the stash (325.6E10), with Pbar stacking (21.6E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8151 at 12:50 AM, put the TeV into shot setup, and established store 8152 at 1:50 AM with an AIL of 300.9E30.

The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/08/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory